# Mitigating Hallucinations in YOLO-based Object Detection Models: A Revisit to Out-of-Distribution Detection

### v1

@@ I INTRODUCTION
  • We conduct extensive experiments that demonstrate the effectiveness of our approach, achieving substantial improvements in both near-OoD and far-OoD detection scenarios.
@@ III-A1 How OoD-only test sets are evaluated
III-A2 Type-1 data error: ID Objects in OoD Test Sets
@@ III-B Issues in Benchmark Datasets - Quantitative Analysis
We quantitatively analyze Type 1 error in the widely used OoD test dataset, OpenImages [22], designed for evaluating OoD detection methods for two ID tasks: PASCAL-VOC [23] and BDD-100K [24]. For brevity, we refer to this subset as OpenImages throughout this paper. We focus on instances where ID objects are erroneously included in the benchmark OoD-only test set, potentially distorting evaluation metrics. We leverage a high-accuracy pre-trained object detection model trained on OpenImagesV7 [22], which recognizes over 500 categories. By running inference on the entire OpenImages dataset, we systematically identify images containing ID objects that should not have been classified as OoD. Our analysis reveals that among the 1,852 images in the dataset, 104 images (5.62%) contain at least one ID object despite being labeled as OoD. Across these images, we detect 220 ID objects, accounting for 12.75% and 13.77% of all detected objects by the models for PASCAL-VOC and BDD-100K tasks (see Table I). To measure the impact of these mislabeling errors on OoD detection performance, we evaluate FPR95 with and without the erroneously labeled images. Our results in Table I show that Type 1 error inflates FPR95 by $9.85\%$ and $5.58\%$ in PASCAL-VOC and BDD-100K tasks, respectively, demonstrating a significant distortion in evaluation metrics. These findings underscore the need for more rigorous dataset curation to ensure the reliability of OoD detection benchmarks.
@@
ID Dataset Type Objects Samples (%)
@@ IV-A Proximal OoD Dataset Construction
To construct a high-quality proximal OoD dataset, we leverage Caltech 256 [25], a diverse object-centric dataset, as a source of candidate samples. Given the large number of categories, manual selection is impractical. Instead, we employ a reasoning-based large language model (LLM) to generate candidate categories that align with our selection principles, followed by manual verification to ensure their validity. Fig. 5 presents the prompt template used to generate candidate proximal OoD categories. Once the proximal OoD categories are identified, we collect approximately 2,000 images and ensure they do not contain any ID objects, as they will serve as background or negative samples during fine-tuning. We systematically analyze the dataset using a pre-trained vision model to enforce this constraint. The model automatically filters out images containing training categories. This automated refinement minimizes ID contamination while preserving the diversity of OoD samples. Finally, the curated dataset is converted into the YOLO format, ensuring seamless integration into the fine-tuning pipeline. This approach streamlines dataset construction and enhances the reliability of proximal OoD selection, enabling more effective mitigation of overconfident predictions.
@@ V-B Fine-tuning
During fine-tuning, we observe the mAP on the ID validation set to analyze model performance change. At the same time, we evaluate the model hallucinations by measuring the number of predictions on OoD data. Fig. 6 illustrate mAP changes over epochs and the trend in hallucination counts on OoD datasets during fine-tuning for the ID task of BDD-100K. The results show that mAP initially declines before gradually improving and returning to its original value by the end of the 20 epochs. This is likely due to the inclusion of proximal OoD data, which presents the model with more challenging backgrounds (similar to ID categories). As it adapts, ID performance temporarily declines. Over time, the model adapts by refining its decision boundaries, regaining ID performance while learning to handle the more challenging OoD data.
Meanwhile, hallucination on both OoD datasets drop sharply in the first epoch and stabilize at a lower level after several epochs (around $20\%$ of the original), suggesting that the model becomes more conservative early in fine-tuning, reducing overconfident predictions. We selected a version for further analysis, reducing OoD detections by up to $80\%$ with only a $1\%$ decrease in ID performance. This version strikes a balance between ID task accuracy and hallucination suppression.
@@ V-C Effectiveness Evaluation
Figure 7: Qualitative comparison of predictions on OoD images using the original and fine-tuned models. (Top) The original model produces hallucinated predictions. (Bottom) The fine-tuned model mitigates hallucinations, enhancing safety in decision-making.
Integrating an OoD Detector To further mitigate hallucinations, we integrate an OoD detector into the fine-tuned model. Our evaluation consists of two steps: first, we assess the performance of various OoD detection methods, and second, we analyze the combined effect of fine-tuning and the best-performing detector. Table V presents the results of our OoD detector evaluation, where BAM and KNN consistently outperform other baselines. Notably, KNN achieves slightly better performance than BAM. Given its superior performance, we use KNN combined with the original and fine-tuned models to examine their respective contributions to reducing hallucinations. Table IV shows that fine-tuning alone significantly improves model reliability, reducing hallucination counts across all datasets. Notably, in the BDD-100K task, the fine-tuned model without an OoD detector significantly reduced hallucinations, with the number of Near-OoD predictions dropping from 297 (Original + BAM) to 145 and Far-OoD predictions decreasing from 290 to 136. When KNN is combined with the fine-tuned model, hallucination counts are further reduced by over $40\%$, demonstrating the additional benefit of leveraging an OoD detector. Specifically, as shown in Table IV, on the BDD-100K benchmark, this combination results in a near $88\%$ reduction in hallucination counts compared to the original model. The original model generated 708 and 641 hallucinations on the near- and far-OoD datasets, respectively, which reduced to 78 and 83 after applying our method. Similarly, for the PASCAL-VOC task, hallucinations decrease by $80.3\%$.

### v2

@@ I INTRODUCTION
  • We conduct extensive experiments that demonstrate the effectiveness of our approach, achieving substantial improvements in both near-OoD and far-OoD detection.
@@ III-A1 How OoD-only test sets are evaluated
III-A2 Type 1 data error: ID Objects in OoD Test Sets
@@ III-B Issues in Benchmark Datasets - Quantitative Analysis
We quantitatively analyze Type 1 error in the widely used OoD test dataset, OpenImages [22], designed for evaluating OoD detection methods for two ID tasks: PASCAL-VOC [23] and BDD-100K [24]. For brevity, we refer to this subset as OpenImages throughout this paper. We focus on instances where ID objects are erroneously included in the benchmark OoD-only test set, potentially distorting evaluation metrics. We leverage a high-accuracy pre-trained object detection model trained on OpenImagesV7 [22], which recognizes over 500 categories. By running inference on the entire OpenImages dataset, we systematically identify images containing ID objects that should not have been classified as OoD. Our analysis reveals that among the 1,852 images in the dataset, 104 images (5.62%) contain at least one ID object despite being labeled as OoD. Across these images, we detect 220 ID objects, accounting for 12.75% and 13.77% of all detected objects by the models for PASCAL-VOC and BDD-100K tasks (see Table I). To measure the impact of these mislabeling errors on OoD detection performance, we evaluate FPR95 with and without the erroneously labeled images. Our results in Table I show that Type 1 error inflates FPR95 by $9.85$ percentage points and $5.58$ percentage points in PASCAL-VOC and BDD-100K tasks, respectively, demonstrating a significant distortion in evaluation metrics. These findings underscore the need for more rigorous dataset curation to ensure the reliability of OoD detection benchmarks.
@@
ID Dataset Type Objects Samples
@@ IV-A Proximal OoD Dataset Construction
To construct a high-quality proximal OoD dataset, we leverage Caltech 256 [25], a diverse object-centric dataset, as a source of candidate samples. Given the large number of categories, manual selection is impractical. Instead, we employ Deepseek-R1, a reasoning-based large language model, to generate candidate categories that align with our selection principles, followed by manual verification to ensure their validity. Fig. 5 presents the prompt template used to generate candidate proximal OoD categories. Once the proximal OoD categories are identified, we collect approximately 2,000 images and ensure they do not contain any ID objects, as they will serve as background or negative samples during fine-tuning. We systematically analyze the dataset using a YOLOv8 model pre-trained on OpenImagesV7 dataset to enforce this constraint. Specifically, we run inference on the collected images and use the FiftyOne library to filter out images containing ID objects based on the inference results. This automated refinement minimizes ID contamination while preserving the diversity of OoD samples. Finally, the curated dataset is converted into the YOLO format, ensuring seamless integration into the fine-tuning pipeline. This approach streamlines dataset construction and enhances the reliability of proximal OoD selection, enabling more effective mitigation of overconfident predictions.
@@ V-B Fine-tuning
During fine-tuning, we observe the mAP on the ID validation set to analyze model performance change. At the same time, we evaluate the model hallucinations by measuring the number of predictions on OoD data. Fig. 6 illustrate mAP changes over epochs and the trend in hallucination counts on OoD datasets during fine-tuning for the ID task of BDD-100K. The results show that mAP initially declines before gradually improving and largely recovering by the end of the 20 epochs, with only a minimal decrease compared to the original baseline. This is likely due to the inclusion of proximal OoD data, which presents the model with more challenging backgrounds (similar to ID categories). As it adapts, ID performance temporarily declines. Over time, the model adapts by refining its decision boundaries, regaining ID performance while learning to handle the more challenging OoD data.
Meanwhile, hallucinations on both OoD datasets drop sharply in the first epoch and stabilize at a lower level after several epochs (around $20\%$ of the original level), suggesting that the model becomes more conservative early in fine-tuning, reducing overconfident predictions. We selected a version for further analysis, reducing OoD detections by up to $80\%$ (relative reduction) with only a $1$ percentage point decrease in ID performance. This version strikes a balance between ID task accuracy and hallucination suppression.
@@ V-C Effectiveness Evaluation
Figure 7: Qualitative comparison of predictions on OoD images using the original and fine-tuned models. (Top) The original model produces hallucinated predictions. (Bottom) The fine-tuned model mitigates them, enhancing decision safety.
Integrating an OoD Detector To further mitigate hallucinations, we integrate an OoD detector into the fine-tuned model. Our evaluation consists of two steps: first, we assess the performance of various OoD detection methods, and second, we analyze the combined effect of fine-tuning and the best-performing detector. Table V presents the results of our OoD detector evaluation, where BAM and KNN consistently outperform other baselines. Notably, KNN achieves slightly better performance than BAM. Given its superior performance, we use KNN combined with the original and fine-tuned models to examine their respective contributions to reducing hallucinations. Table IV shows that fine-tuning alone significantly improves model reliability, reducing hallucination counts across all datasets. Notably, in the BDD-100K task, the fine-tuned model without an OoD detector significantly reduced hallucinations, with the number of Near-OoD predictions dropping from 297 (Original + BAM) to 145 and Far-OoD predictions decreasing from 290 to 136. When KNN is combined with the fine-tuned model, hallucination counts are further reduced by over $40\%$ (relative reduction), demonstrating the additional benefit of leveraging an OoD detector. Specifically, as shown in Table IV, on the BDD-100K benchmark, this combination results in a near $88\%$ relative reduction in hallucination counts compared to the original model. The original model generated 708 and 641 hallucinations on the Near- and Far-OoD datasets, respectively, which reduced to 78 and 83 after applying our method. Similarly, for the PASCAL-VOC task, hallucinations decrease by $80.3\%$ (relative reduction).